Category Archives: Orbitrap status


March 5, 2015

No engineer and no news from the engineer. I have contacted Thermo.


February 27, 2015

The engineer called this afternoon; he is planning the service visit for Wednesday next week.


February 25, 2015

Orbi is down for an unknown time. I have placed a service call this afternoon and will keep you posted. The instrument went down in the middle of Joana’s 4th sample.

All submitted up to this point gel bands have been cleaned-up and are incubating with trypsin. I will lyophilize the peptides and store them in the freezer until the instrument is in operating condition.


February 23, 2015

Shhhh! Orbi is working… for now. Don’t spook it…

My standards ran 2 hrs ago. Gaozhong’s and Josh’s samples are on right now.


February 19, 2015

Orbi is still down, but I think I know why it shuts down. If I am unable to fix it today, I will place a service call. Katsu, Huihun, and Nick – your samples ran last night, and I will send you the data tomorrow. Joana, only first set of your samples is done; second set is ready to go as soon as Orbi is working again. Everyone else – please be patient!


February 13, 2015

Orbi is not ready, it vented for an unknown reason some time last night. I turned it on again this morning, and now we wait and hope that the ghost in the machine is in a better mood today.


February 12, 2015

I can’t believe it’s been 2 weeks since the last post! We’ve been so busy! Today we are recovering from one of our electrical shutdowns: Orbi is baking out, MALDI is ready and is being used by Krista from the Meredith group.

Here’s the current Orbi queue:  Katsu, Nick, Huihun, Joana, Jung (49 hours). MALDI queue: Alex, Jason, Core Facility standards (3 hours).

If Orbi behaves, all analyses should be done by Tuesday next week.


January 30, 2015

What a happy Friday! Orbi has finished Gaozhong’s samples and is running John’s gel bands. After a nice weekend-long wash, Orbi will be running 15 various IP samples (for Scott, Kenny, and Joana).


January 28, 2015

Thanks to the OPP electricians, we have our power restored, and the instruments are pumping down. Orbi will take some time to reach operating conditions, as usual. MALDI is ready for your samples. LCT and QTRAP should be ready tomorrow.

Meanwhile, Scott’s, Joana’s, and John’s gels are incubating with trypsin overnight.